# Is Your Imitation Learning Policy Better than Mine? Policy Comparison with Near-Optimal Stopping

### v1

@@ V-C2 Hardware Evaluation in the Small-Gap Regime
For this task, the initial placement of the carrot is uniformly sampled from three possible locations (left, center, or right) on the counter, and the plate is placed next to the sink (see Figure 9). At the start of each trial, the robot joint angles are initialized such that the gripper is roughly aligned with the carrot initial position. The environment follows a categorical distribution with two outcomes: no object distractors or two object distractors. We consider two environment distributions: Env1 in which there are no distractors with probability $0.8$ and Env2 in which there are no distractors with probability $0.6$. Object distractors are sampled uniformly (without replacement) from the following object categories: orange, apple, green and blue sponges, brown and yellow cubes, eggplant, spoon, and towel. The initial locations of distractors is also sampled uniformly without replacement from four possibilities: on the stove, left of the stove, above the stove, or next to the faucet. We sample 100 environment configurations each for three settings: i) Octo under distribution Env1, ii) OpenVLA under distribution Env1, and iii) Octo under distribution Env2. We then make the following comparisons with STEP: i) Octo (Env1) and Octo (Env2), and ii) Octo (Env1) and OpenVLA (Env1). In the first comparison (Table I, row 5), we test the effect of distribution shift in the probability of distractors being present. Here, the $\hat{p}_{0}$ corresponds to the perturbed distribution Octo (Env2), and $\hat{p}_{1}$ to the nominal distribution Octo (Env1). We find that, while there is an empirical gap (59% vs 68%) at $N=100$, no method returns a significant result at $\alpha^{*}=0.05$. In the second comparison (Table I, row 6), $\hat{p}_{0}$ corresponds to Octo (Env1) and $\hat{p}_{1}$ corresponds to OpenVLA (Env1). We observe no significant gap despite the empirical gap of 8 percentage points in favor of OpenVLA. Importantly, insignificance of the test does not mean that the null hypothesis should be accepted [20]; there remains a possibility that OpenVLA actually outperforms Octo, or that Octo’s performance indeed degrades due to the presence of distractors. However, our budget of $N_{\max}=100$ was likely not sufficient to accumulate enough evidence. In Section IX-C in the Supplement, we further investigate this data insufficiency to show that, if the ground truth values were equal to the empirical success rates (59% vs. 68% and 68% vs. 76%), then we would require $N_{\max}=500$ trials to confidently determine $p_{1}>p_{0}$. This number is an order of magnitude larger than the current norms, reflecting fundamental yet often overlooked challenges in trustworthy policy comparison.
@@ V-C3 Multi-Task Evaluation in SimplerEnv Simulation
Finally, we briefly consider the problem of multi-task and multi-policy extensions to this framework, and illustrate its potential via an example of policy evaluation in simulation (where costs are lower than on hardware, but still can be significant). Concretely, Octo-Small ($\pi_{1}$) and Octo-Base ($\pi_{0}$) [40] are compared in the SimplerEnv [32] simulation environment on three tasks of varying empirical difficulty. On the EggplantInBasket task, the policies each succeed at a near-50% rate, with a gap of 16.4 percentage points. For the SpoonOnTowel task, the gap is larger at 30 percentage points. For the hardest task, StackCube, the performance gap is 3 percentage points. The evaluation statistics are shown in Table I, rows 7-9. Note that the empirical success rates we observed are consistent with the findings of Li et al. [32] that Octo-Small is more performant on these tasks (see their Table V). We seek to evaluate the multitask comparison of the two policies. Letting $p_{s}^{[\tau]}$ denote the performance of Octo-Small and $p_{b}^{[\tau]}$ denote the performance of Octo-Base on task $\tau$, we test:
@@
Many established and sophisticated methods exist to efficiently run multi-hypothesis testing (in this case, we are essentially evaluating three separate hypotheses, one for each task [^5] Note that multi-hypothesis testing can naturally handle the case of multi-policy comparison as well, where we would reduce the test to a set of pairwise comparisons that are examined simultaneously.). As a simple illustration, we use the standard Bonferroni (union bound) correction [15] to evaluate the test: specifically, running each of the individual three tests at level $\alpha^{*}_{[\tau]}=0.01$, we observe the stopping times shown in Table I, rows 7-9 (right-hand side). Via the Bonferroni correction, the combined decision (RejectNull, because every subtest decided RejectNull) expressed in LABEL:Eqn:MultiHypothesis is then confirmed at $\alpha=0.03=\sum_{\tau}\alpha^{*}_{[\tau]}$. As illustrated in Table I, each sequential method saves a substantial number of simulation rollouts on the easiest subtest (SpoonOnTowel). As expected, SAVI begins to struggle when the tests become more challenging, such as in EggplantInBasket. Finally, we observe the weakness of the Lai procedure: in heavily skewed cases, it suffers substantially even compared with SAVI methods, as shown in StackCube. To summarize: naive multitask evaluation requires the aggregation of multiple batches of rollouts, here totaling 1500 per policy (500 per task per policy). Note in rows 7-9 how different the number of requisite trials can be, and therefore how hard it is to reliably run the evaluation using a fixed batch size. On the easiest task, even when tuned to $N_{\text{max}}=500$, the comparison was answered in fewer than 40 rollouts by all sequential methods, a savings of over 90%. On the progressively harder cases the number of required samples increased 5-10 times over the easiest, but our method (STEP) improved substantially over each of the other sequential procedures. In total, STEP would have saved the evaluator an additional 200 rollouts for each of Octo-small and Octo-base for the multitask comparison problem as compared to the current SOTA approaches.
@@ VII Limitations
VIII Conclusion

### v2

@@ V-C2 Hardware Evaluation in the Small-Gap Regime
For this task, the initial placement of the carrot is uniformly sampled from three possible locations (left, center, or right) on the counter, and the plate is placed next to the sink (see Figure 9). At the start of each trial, the robot joint angles are initialized such that the gripper is roughly aligned with the carrot initial position. The environment follows a categorical distribution with two outcomes: no object distractors or two object distractors. We consider two environment distributions: Env1 in which there are no distractors with probability $0.8$ and Env2 in which there are no distractors with probability $0.6$. Object distractors are sampled uniformly (without replacement) from the following object categories: orange, apple, green and blue sponges, brown and yellow cubes, eggplant, spoon, and towel. The initial locations of distractors is also sampled uniformly without replacement from four possibilities: on the stove, left of the stove, above the stove, or next to the faucet. The distractors are physically placed according to a uniform (continuous) distribution within the selected region. We sample 100 environment configurations each for three settings: i) Octo under distribution Env1, ii) OpenVLA under distribution Env1, and iii) Octo under distribution Env2. We then make the following comparisons with STEP: i) Octo (Env1) and Octo (Env2), and ii) Octo (Env1) and OpenVLA (Env1). In the first comparison (Table I, row 7), we test the effect of distribution shift in the probability of distractors being present. Here, the $\hat{p}_{0}$ corresponds to the perturbed distribution Octo (Env2), and $\hat{p}_{1}$ to the nominal distribution Octo (Env1). We find that, while there is an empirical gap (59% vs 68%) at $N=100$, no method returns a significant result at $\alpha^{*}=0.05$. In the second comparison (Table I, row 8), $\hat{p}_{0}$ corresponds to Octo (Env1) and $\hat{p}_{1}$ corresponds to OpenVLA (Env1). We observe no significant gap despite the empirical gap of 8 percentage points in favor of OpenVLA. Importantly, insignificance of the test does not mean that the null hypothesis should be accepted [20]; there remains a possibility that OpenVLA actually outperforms Octo, or that Octo’s performance indeed degrades due to the presence of distractors. However, our budget of $N_{\max}=100$ was likely not sufficient to accumulate enough evidence. In Section X-D in the Supplement, we further investigate this data insufficiency to show that, if the ground truth values were equal to the empirical success rates (59% vs. 68% and 68% vs. 76%), then we would require $N_{\max}=500$ trials to confidently determine $p_{1}>p_{0}$. This number is an order of magnitude larger than the current norms, reflecting fundamental yet often overlooked challenges in trustworthy policy comparison.
@@ V-C3 Multi-Task Evaluation in SimplerEnv Simulation
Finally, we briefly consider the problem of multi-task and multi-policy extensions to this framework, and illustrate its potential via an example of policy evaluation in simulation (where costs are lower than on hardware, but still can be significant). Concretely, Octo-Small ($\pi_{1}$) and Octo-Base ($\pi_{0}$) [40] are compared in the SimplerEnv [32] simulation environment on three tasks of varying empirical difficulty. On the EggplantInBasket task, the policies each succeed at a near-50% rate, with a gap of 16.4 percentage points. For the SpoonOnTowel task, the gap is larger at 30 percentage points. For the hardest task, StackCube, the performance gap is 3 percentage points. The evaluation statistics are shown in Table I, rows 9-11. Note that the empirical success rates we observed are consistent with the findings of Li et al. [32] that Octo-Small is more performant on these tasks (see their Table V). We seek to evaluate the multitask comparison of the two policies. Letting $p_{s}^{[\tau]}$ denote the performance of Octo-Small and $p_{b}^{[\tau]}$ denote the performance of Octo-Base on task $\tau$, we test:
@@
Many established and sophisticated methods exist to efficiently run multi-hypothesis testing (in this case, we are essentially evaluating three separate hypotheses, one for each task [^7] Note that multi-hypothesis testing can naturally handle the case of multi-policy comparison as well, where we would reduce the test to a set of pairwise comparisons that are examined simultaneously.). As a simple illustration, we use the standard Bonferroni (union bound) correction [15] to evaluate the test: specifically, running each of the individual three tests at level $\alpha^{*}_{[\tau]}=0.01$, we observe the stopping times shown in Table I, rows 9-11 (right-hand side). Via the Bonferroni correction, the combined decision (RejectNull, because every subtest decided RejectNull) expressed in LABEL:Eqn:MultiHypothesis is then confirmed at $\alpha=0.03=\sum_{\tau}\alpha^{*}_{[\tau]}$. As illustrated in Table I, each sequential method saves a substantial number of simulation rollouts on the easiest subtest (SpoonOnTowel). As expected, SAVI begins to struggle when the tests become more challenging, such as in EggplantInBasket. Finally, we observe the weakness of the Lai procedure: in heavily skewed cases, it suffers substantially even compared with SAVI methods, as shown in StackCube. To summarize: naive multitask evaluation requires the aggregation of multiple batches of rollouts, here totaling 1500 per policy (500 per task per policy). Note in rows 9-11 how different the number of requisite trials can be, and therefore how hard it is to reliably run the evaluation using a fixed batch size. On the easiest task, even when tuned to $N_{\text{max}}=500$, the comparison was answered in fewer than 40 rollouts by all sequential methods, a savings of over 90%. On the progressively harder cases the number of required samples increased 5-10 times over the easiest, but our method (STEP) improved substantially over each of the other sequential procedures. In total, STEP would have saved the evaluator an additional 200 rollouts for each of Octo-small and Octo-base for the multitask comparison problem as compared to the current SOTA approaches.
@@ VII Limitations
VIII Future Work
There are several concrete extensions of STEP which promise significant benefits to practical robotic evaluation. Most direct is incorporation of more principled multi-policy evaluation, building upon Section V-C3. Additionally, partial credit schema have found increasing importance in long-horizon evaluation [28]; generalizing STEP beyond binary success/failure metrics to discrete (multinoulli) partial credit is direct in theory, although it presents additional challenges in implementation.
IX Conclusion
@@
Acknowledgments
This work was primarily completed during an internship at Toyota Research Institute (TRI). The authors would like to thank Chen Xu for insightful feedback and discussions, as well as the robot teacher team at TRI headquarters for their data collection efforts, in particular Derick Seale and Donovan Jackson. Finally, we thank the anonymous reviewers and the area chair for incisive suggestions for improving the paper.